OpenStudy (anonymous):

Type the equation of the given line in standard form. The line with m = 4 and intersecting 2x -4y = 8 at x = 6.

OpenStudy (anonymous):

HELP ME!!

OpenStudy (anonymous):

What?

OpenStudy (karatechopper):

@hateapex, please do not spam by writing on someone else's question. Thanks.

OpenStudy (anonymous):

find y for x = 6 in 2x -4y = 8 : 2*6-4y=8 4y=4 y=1 so (y-1)=4(x-6) 4x-y=23

OpenStudy (anonymous):

Oh okay. Thanks :)